OpenStudy (anonymous):

I need help understanding functions for math. mainly with problems like f(x) = x^2 - 31 and with the charts for functions. ;~;

OpenStudy (anonymous):

@JayJayTheWereWolfGal1 , @chosenmatt , @Hero

OpenStudy (anonymous):

@Skyz

OpenStudy (alexanderphantomhive):

I am absolutely worthless with math 0_o

OpenStudy (anonymous):

then why are you replying? XD

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I just need help understanding functions is all. just the f(x) and the charts. thats all really

OpenStudy (anonymous):

what do you mean by charts? like a chart of x and y values?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@Alif_Dastegir ?

OpenStudy (anonymous):

oh there you are

OpenStudy (anonymous):

okay basically f(x) = x^2 -31 is the same thing as y = x^2 - 31. So if the value of x is like 10 then it will be f(10) = 10^2 - 31

OpenStudy (anonymous):

Also for a chart for the x and y values you basically plug x value in x^2 - 31 and what ever you get is the y value. For example if you put 1 in for the x value you will get -30 for the y value

OpenStudy (anonymous):

you get it?

OpenStudy (anonymous):

oh!

OpenStudy (anonymous):

ok yeah thank you!

OpenStudy (anonymous):

no problem